OpenStudy (anonymous):

A poster of height 36 in. is mounted on a wall so that its lower edge is 12 in. above the eye level of an observer. How far from the wall should the observer stand so that the viewing angle subtended at his eye by the poster is as large as possible? use optimization to find the length. @Algebra

OpenStudy (dumbcow):

use trig  the goal is to find x which maximize theta \[\theta = \alpha - \beta\] \[\alpha = \tan^{-1} \frac{48}{x}\] \[\beta = \tan^{-1} \frac{12}{x}\] thus \[\theta = \tan^{-1} \frac{48}{x} - \tan^{-1} \frac{12}{x}\] then differentiate and set equal to 0